# Mental Disorders Detection in the Era of Large Language Models

### v1

@@ 4.3 Large Language Models
We conducted experiments with LLMs in various settings. First of all, we evaluated models by 0-shot and 5-shot prompting, considering only normalized probabilities for tokens “0” or “1” in the first generated token, as it was done in MMLU (Hendrycks et al., 2021). We will refer to these settings as “0-shot MMLU” and “5-shot MMLU” correspondingly. Secondly, we employed 0-shot and 5-shot prompting with bigger generation lengths and matched the generated answer to one of the possible classes. Finally, we conducted the finetuning of the models using LoRA (Hu et al., 2022).
We selected a set of relatively small (less than 9B parameters) self-hosted open-source models, either tuned for the Russian language or showing good multilingual capabilities. To the first group belongs SaigaLlama3 8B, a version of Llama 3 8B Instruct (Dubey et al., 2024) tuned on several Russian datasets, as well as models from Vikhr family (Nikolich et al., 2024). We used Vikhr 7B Instruct 0.4, Vikhr 7B Instruct 5.4, and Vikhr Gemma 2B Instruct. The former two are based on Mistral 7B (Jiang et al., 2023) with vocabulary adaptation for the Russian language followed by additional pre-training and instruction tuning. The latter one is based on Gemma2 2B Instruct (Team et al., 2024), additionally trained on Russian data. We also used multilingual models, such as Gemma2 2B Instruct and Gemma2 9B Instruct, as well as Qwen2 7B Instruct (Yang et al., 2024). As we are conducting experiments on sensitive data and with private datasets, we did not consider remotely-hosted models (such as GPT4, Claude) due to the possibility of data leaking.
The full information about used prompts, training hyperparameters, and versions of used models, are presented in Appendix A.
@@ 5.1 Classification Results on Depression and Anxiety Datasets
The data were divided into training and test samples in an 80% by 20% ratio, with stratification by target variable reduced to binary form. All classification reports in this study show results on the test data. The classification report for the best models from each group is presented in Table 2.
@@
| DE | SFT | Linguistic features | 94.00.8 | 95.21.0 | 94.60.8 | 79.34.0 | 75.03.5 | 77.03.3 | 85.82.1 | ||||||||
| SFT | TF-IDF | 91.62.1 | 94.42.2 | 93.01.2 | 74.87.0 | 64.410.3 | 68.56.6 | 80.73.8 | |||||||||
| 5-shot | Vikhr 7B IT 5.4 | 87.060.0 | 82.220.0 | 84.570.0 | 40.740.0 | 50.00.0 | 44.90.0 | 64.730.0 | |||||||||
| 5-shot MMLU | Gemma2 9B IT | 92.960.0 | 73.330.0 | 81.990.0 | 41.460.0 | 77.270.0 | 53.970.0 | 67.980.0 | |||||||||
@@
| SFT | RuBERT | 94.451.16 | 91.111.7 | 92.741.04 | 68.414.04 | 78.034.85 | 72.83.43 | 82.772.21 | |||||||||
| 0-shot | Gemma2 9B IT | 93.330.0 | 62.220.0 | 74.670.0 | 34.620.0 | 81.820.0 | 48.650.0 | 61.660.0 | |||||||||
| DSM | SFT | Linguistic features | 62.81.8 | 59.37.4 | 60.74.0 | 43.72.7 | 47.27.7 | 45.13.8 | 52.92.0 | ||||||||
| SFT | TF-IDF | 62.22.6 | 51.213.9 | 55.38.5 | 42.93.6 | 53.711.4 | 46.94.7 | 51.13.5 | |||||||||
| 5-shot | SaigaLlama3 8B | 100.00.0 | 3.70.0 | 7.140.0 | 40.910.0 | 100.00.0 | 58.060.0 | 32.60.0 | |||||||||
| 5-shot | Vikhr 7B IT 0.4 | 70.970.0 | 81.480.0 | 75.860.0 | 64.290.0 | 50.00.0 | 56.250.0 | 66.060.0 | |||||||||
| 5-shot MMLU | SaigaLlama3 8B | 100.00.0 | 3.70.0 | 7.140.0 | 40.910.0 | 100.00.0 | 58.060.0 | 32.60.0 | |||||||||
| 5-shot MMLU | Vikhr 7B IT 5.4 | 68.750.0 | 81.480.0 | 74.580.0 | 61.540.0 | 44.440.0 | 51.610.0 | 63.090.0 | |||||||||
| LoRA | Qwen2 7B IT | 68.876.39 | 72.228.21 | 70.46.75 | 55.3311.08 | 50.9310.84 | 52.8110.26 | 61.618.32 | |||||||||
| SFT | RuBioRoBERTa | 62.12.98 | 62.966.05 | 62.464.3 | 43.664.43 | 42.594.14 | 43.013.7 | 52.743.67 | |||||||||
| AL | SFT | Linguistic features | 47.421.4 | 36.417.8 | 40.919.2 | 48.52.7 | 68.414.6 | 56.13.7 | 48.58.2 | ||||||||
| SFT | TF-IDF | 61.55.0 | 46.23.1 | 52.62.3 | 51.22.5 | 65.87.9 | 57.54.5 | 55.002.9 | |||||||||
| 5-shot | Gemma2 9B IT | 75.00.0 | 27.270.0 | 40.00.0 | 51.520.0 | 89.470.0 | 65.380.0 | 52.690.0 | |||||||||
| LoRA | Gemma2 2B IT | 60.184.51 | 60.6113.55 | 59.668.17 | 54.877.3 | 53.5110.71 | 53.456.06 | 56.565.4 | |||||||||
| SFT | RuRoBERTa | 56.825.65 | 75.07.76 | 64.495.59 | 52.5410.76 | 33.3312.77 | 40.1712.48 | 52.338.31 | |||||||||
| 0-shot | Gemma2 9B IT | 66.670.0 | 63.640.0 | 65.120.0 | 60.00.0 | 63.160.0 | 61.540.0 | 63.330.0 | |||||||||
| 0-shot MMLU | Gemma2 9B IT | 64.710.0 | 50.00.0 | 56.410.0 | 54.170.0 | 68.420.0 | 60.470.0 | 58.440.0 | |||||||||
| 0-shot MMLU | Gemma2 2B IT | 0.00.0 | 0.00.0 | 0.00.0 | 46.340.0 | 100.00.0 | 63.330.0 | 31.670.0 | |||||||||
| AD | SFT | Linguistic features | 50.83.3 | 43.312.5 | 45.38.1 | 44.72.6 | 51.915.9 | 47.37.4 | 46.31.8 | ||||||||
| SFT | TF-IDF | 54.915.1 | 37.54.8 | 43.22.5 | 44.67.8 | 59.321.0 | 50.412.7 | 46.87.1 | |||||||||
@@
| SFT | RuRoBERTa | 57.632.52 | 56.674.71 | 57.073.19 | 52.82.82 | 53.74.14 | 53.172.84 | 55.122.58 | |||||||||
| 0-shot | Vikhr 7B IT 5.4 | 83.330.0 | 25.00.0 | 38.460.0 | 53.120.0 | 94.440.0 | 68.00.0 | 53.230.0 | |||||||||
| 0-shot | Qwen2 7B IT | 69.570.0 | 80.00.0 | 74.420.0 | 73.330.0 | 61.110.0 | 66.670.0 | 70.540.0 | |||||||||
| 0-shot MMLU | Qwen2 7B IT | 73.680.0 | 70.00.0 | 71.790.0 | 68.420.0 | 72.220.0 | 70.270.0 | 71.030.0 | |||||||||
| AC | SFT | Linguistic features | 64.816.0 | 45.621.0 | 47.019.6 | 48.73.6 | 60.520.1 | 52.67.3 | 49.87.8 | ||||||||
| SFT | TF-IDF | 56.32.1 | 63.34.2 | 59.52.5 | 48.93.3 | 41.75.6 | 44.94.3 | 52.22.7 | |||||||||
| 5-shot | Gemma2 9B IT | 100.00.0 | 15.560.0 | 26.920.0 | 50.00.0 | 100.00.0 | 66.670.0 | 46.790.0 | |||||||||
| 5-shot MMLU | SaigaLlama3 8B | 100.00.0 | 13.330.0 | 23.530.0 | 49.350.0 | 100.00.0 | 66.090.0 | 44.810.0 | |||||||||
| LoRA | Vikhr 7B IT 5.4 | 61.872.07 | 62.968.18 | 62.134.3 | 55.53.03 | 53.957.25 | 54.333.58 | 58.232.38 | |||||||||
@@
| 0-shot | SaigaLlama3 8B | 67.740.0 | 46.670.0 | 55.260.0 | 53.850.0 | 73.680.0 | 62.220.0 | 58.740.0 | |||||||||
| 0-shot MMLU | SaigaLlama3 8B | 70.370.0 | 42.220.0 | 52.780.0 | 53.570.0 | 78.950.0 | 63.830.0 | 58.30.0 |
@@ 5.1.1 DE Dataset
The best scores of the F1 for the pathology class and F1-macro in this experiment are achieved on the essay dataset (DE) by the finetuned LLM model with 88.4% F1-macro and 81.1% F1 score for the pathology class. The model trained on the linguistic features shows similar results with 85.8% F1-macro and 77.0% F1-pathology.
In general, the three best results on the DE dataset were achieved by finetuned models, which can be linked to the bigger dataset size and to the longer texts in the dataset, which, in turn, are crucial for supervised finetuning. For all other datasets methods without finetuning significantly outperform SFT and LoRA.
@@
| D-all | 5-shot MMLU | Gemma2 9B IT | 89.330.0 | 97.550.0 | 93.260.0 | 69.230.0 | 32.140.0 | 43.90.0 | 68.580.0 | ||||||||
| LoRA | Gemma2 2B IT | 92.330.94 | 97.141.04 | 94.670.72 | 76.446.36 | 52.986.33 | 62.335.6 | 78.53.13 | |||||||||
| SFT | RuBERT | 92.260.52 | 96.221.56 | 94.190.68 | 72.1210.1 | 52.983.81 | 60.62.87 | 77.391.72 | |||||||||
| 0-shot | Gemma2 9B IT | 93.280.0 | 76.690.0 | 84.180.0 | 33.330.0 | 67.860.0 | 44.710.0 | 64.440.0 | |||||||||
| SFT | Linguistic features | 88.10.6 | 88.13.4 | 88.01.6 | 51.65.8 | 50.74.1 | 50.82.6 | 69.41.9 | |||||||||
| SFT | TF-IDF | 89.51.0 | 84.21.0 | 86.80.6 | 47.51.8 | 59.14.2 | 52.62.5 | 69.71.5 | |||||||||
| A-all | 5-shot | Gemma2 2B IT | 72.00.0 | 20.690.0 | 32.140.0 | 49.640.0 | 90.670.0 | 64.150.0 | 48.150.0 | ||||||||
@@
| SFT | RuRoBERTa | 61.533.69 | 60.1513.8 | 59.684.47 | 54.390.97 | 54.4418.08 | 52.2711.87 | 55.973.88 | |||||||||
| SFT | BERT | 59.441.74 | 46.933.66 | 52.332.1 | 50.430.98 | 62.674.81 | 55.822.21 | 54.081.08 | |||||||||
| 0-shot | Qwen2 7B IT | 60.820.0 | 67.820.0 | 64.130.0 | 56.920.0 | 49.330.0 | 52.860.0 | 58.490.0 | |||||||||
| 0-shot MMLU | SaigaLlama3 8B | 69.770.0 | 34.480.0 | 46.150.0 | 52.10.0 | 82.670.0 | 63.920.0 | 55.040.0 | |||||||||
| 0-shot MMLU | Qwen2 7B IT | 62.790.0 | 62.070.0 | 62.430.0 | 56.580.0 | 57.330.0 | 56.950.0 | 59.690.0 | |||||||||
| SFT | Linguistic features | 53.42.9 | 39.719.3 | 42.017.9 | 46.12.4 | 60.220.4 | 51.07.8 | 46.56.2 | |||||||||
| SFT | TF-IDF | 51.80.7 | 40.015.6 | 43.310.8 | 44.32.1 | 56.417.8 | 48.68.5 | 46.02.4 |
@@
| DE to DSM test | LoRA | SaigaLlama3 8B | 58.493.68 | 83.9513.97 | 68.737.39 | 40.3413.85 | 12.044.99 | 17.215.85 | 42.974.2 | ||||||||
@@
| LoRA | Vikhr 7B IT 5.4 | 58.245.42 | 67.2821.75 | 60.2411.51 | 31.337.8 | 25.9324.36 | 25.0915.4 | 42.664.75 | |||||||||
| LoRA | Qwen2 7B IT | 61.143.36 | 61.1118.61 | 59.7511.42 | 43.628.81 | 42.5915.93 | 41.258.0 | 50.55.32 | |||||||||
| LoRA | VikhrGemma 2B IT | 60.842.46 | 71.616.93 | 64.98.82 | 42.547.31 | 31.4815.27 | 33.9712.07 | 49.445.25 | |||||||||
| SFT | RuRoBERTa | 60.967.01 | 41.9826.59 | 46.0714.36 | 33.7515.28 | 59.2627.34 | 42.9519.5 | 44.515.07 | |||||||||
| SFT | RuBioRoBERTa | 61.652.51 | 50.023.88 | 52.5412.11 | 34.5415.55 | 52.7824.59 | 41.5718.71 | 47.055.29 | |||||||||
| SFT | RuBERT | 68.010.29 | 27.789.01 | 38.6810.16 | 42.662.8 | 80.569.49 | 55.644.19 | 47.165.8 | |||||||||

### v3

@@ 4.3 Large Language Models
We conducted experiments with LLMs in various settings. First of all, we evaluated models by 0-shot and 5-shot prompting, considering only normalized probabilities for tokens “0” or “1” in the first generated token, as it was done in MMLU (Hendrycks et al., 2021). We will refer to these settings as “0-shot MMLU” and “5-shot MMLU” correspondingly. Secondly, we employed 0-shot and 5-shot prompting with bigger generation lengths and matched the generated answer to one of the possible classes with string matching. Finally, we conducted the fine-tuning of the models using LoRA (Hu et al., 2022).
We selected a set of relatively small (less than 9B parameters) self-hosted open-source models, either fine-tuned for the Russian language or showing good multilingual capabilities. To the first group belongs SaigaLlama3 8B, a version of Llama 3 8B Instruct (Grattafiori et al., 2024) fine-tuned on several Russian datasets, as well as models from the Vikhr family (Nikolich et al., 2024). We used Vikhr 7B Instruct 0.4, Vikhr 7B Instruct 5.4, and Vikhr Gemma 2B Instruct. The former two are based on Mistral 7B (Jiang et al., 2023) with vocabulary adaptation for the Russian language, followed by additional pretraining and instruction tuning. The latter one is based on Gemma2 2B Instruct (Team et al., 2024), additionally trained on Russian data. We also used multilingual models, such as Gemma2 2B Instruct and Gemma2 9B Instruct, as well as Qwen2 7B Instruct (Yang et al., 2024). As we are conducting experiments on sensitive data and with private datasets, we did not consider remotely-hosted models (such as GPT-4, Claude) due to the possibility of data leakage.
The full information about used prompts, training hyperparameters, and versions of used models is presented in Appendix˜A.
@@ 5.1 Classification Results on Depression and Anxiety Datasets
The data were divided into training and test samples in an 80% by 20% ratio, with stratification by the target variable reduced to binary form. All classification reports in this study show results on the test data. The classification report for the best models from each group is presented in Table˜2.
@@
| DE | SFT | Linguistic features | 94.00 0.80 | 95.20 1.00 | 94.60 0.80 | 79.30 4.00 | 75.00 3.50 | 77.00 3.30 | 85.80 2.10 | ||||||||
| SFT | TF-IDF | 91.60 2.10 | 94.40 2.20 | 93.00 1.20 | 74.80 7.00 | 64.40 10.30 | 68.50 6.60 | 80.70 3.80 | |||||||||
| 5-shot | Vikhr 7B IT 5.4 | 87.06 0.00 | 82.22 0.00 | 84.57 0.00 | 40.74 0.00 | 50.00 0.00 | 44.90 0.00 | 64.73 0.00 | |||||||||
| 5-shot MMLU | Gemma2 9B IT | 93.11 0.08 | 75.11 0.89 | 83.15 0.58 | 43.16 0.85 | 77.27 0.00 | 55.38 0.71 | 69.26 0.64 | |||||||||
@@
| SFT | RuBERT | 94.45 1.16 | 91.11 1.70 | 92.74 1.04 | 68.41 4.04 | 78.03 4.85 | 72.80 3.43 | 82.77 2.21 | |||||||||
| 0-shot | Gemma2 9B IT | 93.33 0.00 | 62.22 0.00 | 74.67 0.00 | 34.62 0.00 | 81.82 0.00 | 48.65 0.00 | 61.66 0.00 | |||||||||
| DSM | SFT | Linguistic features | 62.80 1.80 | 59.30 7.40 | 60.70 4.00 | 43.70 2.70 | 47.20 7.70 | 45.10 3.80 | 52.90 2.00 | ||||||||
| SFT | TF-IDF | 62.20 2.60 | 51.20 13.90 | 55.30 8.50 | 42.90 3.60 | 53.70 11.40 | 46.90 4.70 | 51.10 3.50 | |||||||||
| 5-shot | SaigaLlama3 8B | 100.00 0.00 | 3.70 0.00 | 7.14 0.00 | 40.91 0.00 | 100.00 0.00 | 58.06 0.00 | 32.60 0.00 | |||||||||
| 5-shot | Vikhr 7B IT 0.4 | 69.19 0.89 | 81.48 0.00 | 74.83 0.51 | 62.09 1.10 | 45.56 2.22 | 52.54 1.85 | 63.69 1.18 | |||||||||
| 5-shot MMLU | SaigaLlama3 8B | 100.00 0.00 | 3.70 0.00 | 7.14 0.00 | 40.91 0.00 | 100.00 0.00 | 58.06 0.00 | 32.60 0.00 | |||||||||
| 5-shot MMLU | Vikhr 7B IT 5.4 | 68.75 0.00 | 81.48 0.00 | 74.58 0.00 | 61.54 0.00 | 44.44 0.00 | 51.61 0.00 | 63.09 0.00 | |||||||||
| LoRA | Qwen2 7B IT | 68.87 6.39 | 72.22 8.21 | 70.40 6.75 | 55.33 11.08 | 50.93 10.84 | 52.81 10.26 | 61.61 8.32 | |||||||||
| SFT | RuBioRoBERTa | 62.10 2.98 | 62.96 6.05 | 62.46 4.30 | 43.66 4.43 | 42.59 4.14 | 43.01 3.70 | 52.74 3.67 | |||||||||
| AL | SFT | Linguistic features | 47.40 21.40 | 36.40 17.80 | 40.90 19.20 | 48.50 2.70 | 68.40 14.60 | 56.10 3.70 | 48.50 8.20 | ||||||||
| SFT | TF-IDF | 61.50 5.00 | 46.20 3.10 | 52.60 2.30 | 51.20 2.50 | 65.80 7.90 | 57.50 4.50 | 55.00 2.90 | |||||||||
| 5-shot | Gemma2 9B IT | 75.00 0.00 | 27.27 0.00 | 40.00 0.00 | 51.52 0.00 | 89.47 0.00 | 65.38 0.00 | 52.69 0.00 | |||||||||
| LoRA | Gemma2 2B IT | 60.18 4.51 | 60.61 13.55 | 59.66 8.17 | 54.87 7.30 | 53.51 10.71 | 53.45 6.06 | 56.56 5.40 | |||||||||
| SFT | RuRoBERTa | 56.82 5.65 | 75.00 7.76 | 64.49 5.59 | 52.54 10.76 | 33.33 12.77 | 40.17 12.48 | 52.33 8.31 | |||||||||
| 0-shot | Gemma2 9B IT | 66.67 0.00 | 63.64 0.00 | 65.12 0.00 | 60.00 0.00 | 63.16 0.00 | 61.54 0.00 | 63.33 0.00 | |||||||||
| 0-shot MMLU | Gemma2 2B IT | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 46.34 0.00 | 100.00 0.00 | 63.33 0.00 | 31.67 0.00 | |||||||||
| 0-shot MMLU | Qwen2 7B IT | 61.03 0.52 | 85.45 1.82 | 71.20 0.99 | 68.73 2.55 | 36.84 0.00 | 47.95 0.64 | 59.58 0.82 | |||||||||
| AD | SFT | Linguistic features | 50.80 3.30 | 43.30 12.50 | 45.30 8.10 | 44.70 2.60 | 51.90 15.90 | 47.30 7.40 | 46.30 1.80 | ||||||||
| SFT | TF-IDF | 54.90 15.10 | 37.50 4.80 | 43.20 2.50 | 44.60 7.80 | 59.30 21.00 | 50.40 12.70 | 46.80 7.10 | |||||||||
@@
| SFT | RuRoBERTa | 57.63 2.52 | 56.67 4.71 | 57.07 3.19 | 52.80 2.82 | 53.70 4.14 | 53.17 2.84 | 55.12 2.58 | |||||||||
| 0-shot | Vikhr 7B IT 5.4 | 83.33 0.00 | 25.00 0.00 | 38.46 0.00 | 53.12 0.00 | 94.44 0.00 | 68.00 0.00 | 53.23 0.00 | |||||||||
| 0-shot | Qwen2 7B IT | 67.25 1.16 | 80.00 0.00 | 73.07 0.68 | 71.81 0.76 | 56.67 2.22 | 63.33 1.67 | 68.20 1.17 | |||||||||
| 0-shot MMLU | Vikhr 7B IT 5.4 | 83.33 0.00 | 25.00 0.00 | 38.46 0.00 | 53.12 0.00 | 94.44 0.00 | 68.00 0.00 | 53.23 0.00 | |||||||||
| 0-shot MMLU | Qwen2 7B IT | 70.74 1.47 | 70.00 0.00 | 70.36 0.72 | 67.02 0.70 | 67.78 2.22 | 67.39 1.44 | 68.87 1.08 | |||||||||
| AC | SFT | Linguistic features | 64.80 16.00 | 45.60 21.00 | 47.00 19.60 | 48.70 3.60 | 60.50 20.10 | 52.60 7.30 | 49.80 7.80 | ||||||||
| SFT | TF-IDF | 56.30 2.10 | 63.30 4.20 | 59.50 2.50 | 48.90 3.30 | 41.70 5.60 | 44.90 4.30 | 52.20 2.70 | |||||||||
| 5-shot | Qwen2 7B IT | 84.89 3.56 | 26.22 5.33 | 39.85 6.96 | 52.08 1.72 | 94.74 0.00 | 67.20 1.46 | 53.52 4.21 | |||||||||
| 5-shot MMLU | SaigaLlama3 8B | 100.00 0.00 | 13.33 0.00 | 23.53 0.00 | 49.35 0.00 | 100.00 0.00 | 66.09 0.00 | 44.81 0.00 | |||||||||
| LoRA | Vikhr 7B IT 5.4 | 61.87 2.07 | 62.96 8.18 | 62.13 4.30 | 55.50 3.03 | 53.95 7.25 | 54.33 3.58 | 58.23 2.38 | |||||||||
@@
| 0-shot | SaigaLlama3 8B | 66.05 0.85 | 46.67 0.00 | 54.69 0.29 | 53.12 0.36 | 71.58 1.05 | 60.98 0.62 | 57.84 0.45 | |||||||||
| 0-shot MMLU | Vikhr 7B IT 0.4 | 69.15 2.43 | 41.78 0.89 | 52.09 1.39 | 53.04 1.06 | 77.89 2.11 | 63.11 1.45 | 57.60 1.42 |
@@ 5.1.1 DE Dataset
The best scores of the F1 for the pathology class and F1-macro in this experiment are achieved on the essay dataset (DE) by the fine-tuned LLM model with 88.4% F1-macro and 81.1% F1 score for the pathology class. The model trained on the linguistic features shows comparable results with 85.8% F1-macro and 77.0% F1-pathology.
In general, the three best results on the DE dataset were achieved by fine-tuned models, which can be linked to the bigger dataset size and to the longer texts in the dataset, which, in turn, are crucial for supervised fine-tuning. For all other datasets, various prompting methods without finetuning significantly outperform SFT and LoRA.
@@
| D-all | 5-shot MMLU | Gemma2 9B IT | 88.83 0.25 | 96.56 0.49 | 92.53 0.36 | 59.56 4.84 | 29.29 1.43 | 39.26 2.32 | 65.90 1.34 | ||||||||
| LoRA | Gemma2 2B IT | 92.33 0.94 | 97.14 1.04 | 94.67 0.72 | 76.44 6.36 | 52.98 6.33 | 62.33 5.60 | 78.50 3.13 | |||||||||
| SFT | RuBERT | 92.26 0.52 | 96.22 1.56 | 94.19 0.68 | 72.12 10.10 | 52.98 3.81 | 60.60 2.87 | 77.39 1.72 | |||||||||
| 0-shot | Gemma2 9B IT | 93.28 0.00 | 76.69 0.00 | 84.18 0.00 | 33.33 0.00 | 67.86 0.00 | 44.71 0.00 | 64.44 0.00 | |||||||||
| 0-shot MMLU | Gemma2 9B IT | 94.87 0.00 | 68.10 0.00 | 79.29 0.00 | 29.73 0.00 | 78.57 0.00 | 43.14 0.00 | 61.21 0.00 | |||||||||
| SFT | Linguistic features | 88.10 0.60 | 88.10 3.40 | 88.00 1.60 | 51.60 5.80 | 50.70 4.10 | 50.80 2.60 | 69.40 1.90 | |||||||||
| SFT | TF-IDF | 89.50 1.00 | 84.20 1.00 | 86.80 0.60 | 47.50 1.80 | 59.10 4.20 | 52.60 2.50 | 69.70 1.50 | |||||||||
| A-all | 5-shot MMLU | SaigaLlama3 8B | 80.51 5.64 | 11.26 0.46 | 19.76 0.88 | 48.46 0.41 | 96.80 1.07 | 64.59 0.60 | 42.18 0.74 | ||||||||
@@
| SFT | RuRoBERTa | 61.53 3.69 | 60.15 13.80 | 59.68 4.47 | 54.39 0.97 | 54.44 18.08 | 52.27 11.87 | 55.97 3.88 | |||||||||
| SFT | BERT | 59.44 1.74 | 46.93 3.66 | 52.33 2.10 | 50.43 0.98 | 62.67 4.81 | 55.82 2.21 | 54.08 1.08 | |||||||||
| 0-shot | Qwen2 7B IT | 60.01 0.41 | 69.66 0.92 | 64.47 0.17 | 56.72 0.10 | 46.13 1.60 | 50.87 0.99 | 57.67 0.41 | |||||||||
| 0-shot | Gemma2 2B IT | 71.44 0.87 | 26.44 0.00 | 38.59 0.13 | 50.69 0.15 | 87.73 0.53 | 64.26 0.27 | 51.42 0.20 | |||||||||
| 0-shot MMLU | Qwen2 7B IT | 62.67 0.06 | 64.83 1.38 | 63.72 0.65 | 57.51 0.47 | 55.20 1.07 | 56.32 0.32 | 60.02 0.17 | |||||||||
| SFT | Linguistic features | 53.40 2.90 | 39.70 19.30 | 42.00 17.90 | 46.10 2.40 | 60.20 20.40 | 51.00 7.80 | 46.50 6.20 | |||||||||
| SFT | TF-IDF | 51.80 0.70 | 40.00 15.60 | 43.30 10.80 | 44.30 2.10 | 56.40 17.80 | 48.60 8.50 | 46.00 2.40 |
@@
| DE to DSM test | LoRA | SaigaLlama3 8B | 58.49 3.68 | 83.95 13.97 | 68.73 7.39 | 40.34 13.85 | 12.04 4.99 | 17.21 5.85 | 42.97 4.20 | ||||||||
@@
| LoRA | Vikhr 7B IT 5.4 | 58.24 5.42 | 67.28 21.75 | 60.24 11.51 | 31.33 7.80 | 25.93 24.36 | 25.09 15.40 | 42.66 4.75 | |||||||||
| LoRA | Qwen2 7B IT | 61.14 3.36 | 61.11 18.61 | 59.75 11.42 | 43.62 8.81 | 42.59 15.93 | 41.25 8.00 | 50.50 5.32 | |||||||||
| LoRA | VikhrGemma 2B IT | 60.84 2.46 | 71.60 16.93 | 64.90 8.82 | 42.54 7.31 | 31.48 15.27 | 33.97 12.07 | 49.44 5.25 | |||||||||
| SFT | RuRoBERTa | 60.96 7.01 | 41.98 26.59 | 46.07 14.36 | 33.75 15.28 | 59.26 27.34 | 42.95 19.50 | 44.51 5.07 | |||||||||
| SFT | RuBioRoBERTa | 61.65 2.51 | 50.00 23.88 | 52.54 12.11 | 34.54 15.55 | 52.78 24.59 | 41.57 18.71 | 47.05 5.29 | |||||||||
| SFT | RuBERT | 68.00 10.29 | 27.78 9.01 | 38.68 10.16 | 42.66 2.80 | 80.56 9.49 | 55.64 4.19 | 47.16 5.80 | |||||||||